One side of a triangle is 2 cm
shorter than the base. The other side is 3 cm longer than
the base. What lengths of the base will allow the
perimeter to be greater than 19 cm?”

Answer by london maths tutor(243)

Let's base lenght be "b"
Shorter side = (b - 2)cm
Longer side = (b + 3)cm
Perimeter is the sum of all lengths:
b + b - 2 + b + 3 = 3b - 1
3b - 1 > 19
3b > 19 + 1
3b > 20 cm
b > (20÷3) cm
b > 6.666 cm
Answer: The base length has to be greater than 6.67cm.